# On the Power Saving in High-Speed Ethernet-based Networks for Supercomputers and Data Centers

### v1

@@ 4.4 Evaluation for the AlexNet application
Our AlexNet traffic pattern is extracted from a training session, and it is shown in Fig. 15. During this time, the main collective communications happen mainly during the back-propagation after each layer has been trained, and gradients are averaged by means of an AllReduce operation. We can see that the network is mostly idling, but there are periodic traffic bursts due to the collective communications. During the computation-heavy periods, however, we have opportunities for power-saving. Without any power-saving applied, the system consumed 397.437 MJ in total, of which 323.786 MJ (81.469 %) are consumed by the nodes, and the rest comes from the network.
@@ 4.4.1 Results with fixed ${t}_{PDT}$ values
Fig. 16(a) shows the effect on execution time caused by employing every $t_{PDT}$ value. As we can see, the overhead caused is similar to values above the nanosecond range, in which 90 % of inactivity periods fit, as shown in Fig. 1(d). Because that is also the $t_{w}$ for Fast Wake, we find packet coalescence on output buffers on their way to their destinations, compensating for the overhead. This is why we do not see a steep increase in execution time when very restrictive $t_{PDT}$ values are employed.
@@
Another hint of packet coalescence occurring when using Fast Wake can be seen in Fig. 16(b). We can see that the latency increase is minimal compared to Deep Sleep.
The benefit of using Deep Sleep instead of Fast Wake is the tradeoff of energy savings for performance. As we can see in Fig. 16(c), using Deep Sleep leads to more energy savings. The increase in energy saving as the $t_{PDT}$ increases is indicative that the potential for energy saving is hindered by the high number of transitions that occur. This also explains why using a $t_{PDT}$ of 1 second reverts this trend, because then it is unused ports with active PDT timers that prevent energy saving.
@@ 4.4.2 Results using PerfBound and PerfBoundCorrect
As we can see in Fig. 17(a), the results show the same tendencies as with the previous applications. However, this time we manage not only the same reduction in execution time overhead, but we also get that overhead to be below the performance degradation threshold using PerfBoundCorrect.
@@
For packet latency, the results are shown in different figures for Fast Wake and Deep Sleep due to the difference in the obtained results. We can see in Fig. 17(c) that even if latency values are worse for PerfBoundCorrect than for PerfBound, the increase is still negligible when the packet latency is within the microsecond scale. Using Deep Sleep, even if the increase is small, we still manage to improve upon PerfBound in all cases.
@@ Acknowledgements
This work has been supported by the Junta de Comunidades de Castilla-La Mancha under projects SBPLY/21/180225/000103 and SBPLY/21/180501/000248, and by the Spanish Ministry of Science and Innovation (MCIN) / Agencia Estatal de Investigación (AEI) under project PID2021-123627OB-C52, co-financed by the European Regional Development Fund (FEDER). We also acknowledge support from the PERTE-Chip grants (UCLM Chair, TSI-069100-2023-0014) from the Spanish Ministry of Digital Transformation and Public Service.

### v2

@@ 4.4 Evaluation for the AlexNet application
Our AlexNet traffic pattern is extracted from a training session, and it is shown in Figure 15. During this time, the main collective communications happen mainly during the back-propagation after each layer has been trained, and gradients are averaged by means of an AllReduce operation. We can see that the network is mostly idling, but there are periodic traffic bursts due to the collective communications. During the computation-heavy periods, however, we have opportunities for power-saving. Without any power-saving applied, the system consumed 397.437 MJ in total, of which 323.786 MJ (81.469 %) are consumed by the nodes, and the rest comes from the network.
@@ 4.4.1 Results with fixed ${t}_{PDT}$ values
Figure 16(a) shows the effect on execution time caused by employing every $t_{PDT}$ value. As we can see, the overhead caused is similar to values above the nanosecond range, in which 90 % of inactivity periods fit, as shown in Figure 1(d). Because that is also the $t_{w}$ for Fast Wake, we find packet coalescence on output buffers on their way to their destinations, compensating for the overhead. This is why we do not see a steep increase in execution time when very restrictive $t_{PDT}$ values are employed.
@@
Another hint of packet coalescence occurring when using Fast Wake can be seen in Figure 16(b). We can see that the latency increase is minimal compared to Deep Sleep.
The benefit of using Deep Sleep instead of Fast Wake is the tradeoff of energy savings for performance. As we can see in Figure 16(c), using Deep Sleep leads to more energy savings. The increase in energy saving as the $t_{PDT}$ increases is indicative that the potential for energy saving is hindered by the high number of transitions that occur. This also explains why using a $t_{PDT}$ of 1 second reverts this trend, because then it is unused ports with active PDT timers that prevent energy saving.
@@ 4.4.2 Results using PerfBound and PerfBoundCorrect
As we can see in Figure 17(a), the results show the same tendencies as with the previous applications. However, this time we manage not only the same reduction in execution time overhead, but we also get that overhead to be below the performance degradation threshold using PerfBoundCorrect.
@@
For packet latency, the results are shown in different figures for Fast Wake and Deep Sleep due to the difference in the obtained results. We can see in Figure 17(c) that even if latency values are worse for PerfBoundCorrect than for PerfBound, the increase is still negligible when the packet latency is within the microsecond scale. Using Deep Sleep, even if the increase is small, we still manage to improve upon PerfBound in all cases.
@@ Acknowledgements
This work has been supported by the Junta de Comunidades de Castilla-La Mancha under projects SBPLY/21/180225/000103, TED2021-130233B-C31 and SBPLY/21/180501/000248, and by the Spanish Ministry of Science and Innovation (MCIN) / Agencia Estatal de Investigación (AEI) under project PID2021-123627OB-C52, co-financed by the European Regional Development Fund (FEDER). We also acknowledge support from the PERTE-Chip grants (UCLM Chair, TSI-069100-2023-0014) from the Spanish Ministry of Digital Transformation and Public Service.